# Matrix Manifold Neural Networks++

### v1

@@ 3.3 MLR in Structure Spaces
Motivated by the works in Nguyen (2022a); Nguyen & Yang (2023), in this section, we aim to build MLR on SPSD manifolds. For any $\mathbf{P}\in\operatorname{S}^{+}_{n,p}$, we consider the decomposition $\mathbf{P}=\mathbf{U}_{P}\mathbf{S}_{P}\mathbf{U}^{T}_{P}$, where $\mathbf{U}_{P}\in\widetilde{\operatorname{Gr}}_{n,p}$ and $\mathbf{S}_{P}\in\operatorname{Sym}_{p}^{+}$. Each element of $\operatorname{S}^{+}_{n,p}$ can be seen as a flat $p$-dimensional ellipsoid in $\mathbb{R}^{n}$ (Bonnabel et al., 2013). The flat ellipsoid belongs to a $p$-dimensional subspace spanned by the columns of $\mathbf{U}_{P}$, while the $p\times p$ SPD matrix $\mathbf{S}_{P}$ defines the shape of the ellipsoid in $\operatorname{Sym}_{p}^{+}$. A canonical representation of $\mathbf{P}$ in structure space $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+}$ is computed by identifying a common subspace and then rotating $\mathbf{U}_{P}$ to this subspace. The SPD matrix $\mathbf{S}_{P}$ is rotated accordingly to reflect the changes of $\mathbf{U}_{P}$. Details of these computations are given in Appendix H.
@@
where ~$\widetilde{\oplus}_{gr}$ is the binary operation in $\widetilde{\operatorname{Gr}}_{n,p}$ (see Appendix G.6 for the definition of ~$\widetilde{\oplus}_{gr}$).
@@
where ~$\widetilde{\ominus}_{gr}$ is the inverse operation in $\widetilde{\operatorname{Gr}}_{n,p}$ (see Appendix G.6 for the definition of ~$\widetilde{\ominus}_{gr}$).
Our construction of the binary and inverse operations in $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+,g}$ is clearly advantageous compared to the method in Nguyen (2022a) since this method does not preserve the information about the subspaces of the terms involved in these operations. In addition to the binary and inverse operations, we also need to define the inner product in structure spaces.
@@
where $\lambda>0$, .,.$\langle.,.\rangle^{gr}$ is the Grassmann inner product (Nguyen & Yang, 2023) (see Appendix G.7).
The key idea to generalize MLR to a Riemannian manifold is to change the margin to reflect the geometry of the considered manifold (a formulation of MLR from the perspective of distances to hyperplanes is given in Appendix C). This requires the notions of hyperplanes and margin in the considered manifold that are referred to as hypergyroplanes and pseudo-gyrodistances (Nguyen & Yang, 2023), respectively. In our case, the definition of hypergyroplanes in structure spaces, suggested by Proposition 3.2, can be given below.
@@
where .$\|.\|^{gr}$ and .$\|.\|^{g}$ are the norms induced by the Grassmann and SPD inner products, respectively.
@@ 3.4.1 Grassmann Logarithmic Map in The Projector Perspective
The Grassmann logarithmic map is given (Batzies et al., 2015; Bendokat et al., 2020) by
@@ 3.4.2 Graph Convolutional Networks on Grassmann Manifolds
We propose to extend GCNs to Grassmann geometry using an approach similar to Chami et al. (2019); Zhao et al. (2023). Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, $\mathbf{x}_{i}^{l},i\in\mathcal{V}$ be the embedding of node $i$ at layer $l$ ($l=0$ indicates input node features), $\mathcal{N}(i)=\{j:(i,j)\in\mathcal{E}\}$ be the set of neighbors of $i\in\mathcal{V}$, $\mathbf{W}^{l}$ and $\mathbf{b}^{l}$ be the weight and bias for layer $l$, and $\sigma(.)$ be a non-linear activation function. A basic GCN message-passing update (Zhao et al., 2023) can be expressed as
@@
For the aggregation operation, the weights $w_{ij}$ can be computed using different methods (Kipf & Welling, 2017; Hamilton et al., 2017). Let $\mathbf{X}_{i}^{l}\in\operatorname{Gr}_{n,p},i\in\mathcal{V}$ be the Grassmann embedding of node $i$ at layer $l$. For feature transformation on Grassmann manifolds, we use isometry maps based on left Grassmann gyrotranslations (Nguyen & Yang, 2023), i.e.,
@@
where $\mathbf{P}_{i}^{l}$ and $\mathbf{Q}_{i}^{l}$ are the input and output node features of the aggregation operation, and $k_{i,j}=|\mathcal{N}(i)|^{-\frac{1}{2}}|\mathcal{N}(j)|^{-\frac{1}{2}}$ represents the relative importance of node $j$ to node $i$. For any $\mathbf{X}\in\operatorname{Gr}_{n,p}$, let $\exp([\operatorname{Log}^{gr}_{\mathbf{I}_{n,p}}(\mathbf{X}),\mathbf{I}_{n,p}])\widetilde{\mathbf{I}}_{n,p}=\mathbf{V}\mathbf{U}$ be a QR decomposition, where $\widetilde{\mathbf{I}}_{n,p}=\begin{bmatrix}\mathbf{I}_{p}\\ 0\end{bmatrix}\in\operatorname{M}_{n,p}$, $\mathbf{V}\in\operatorname{M}_{n,p}$ is a matrix with orthonormal columns, and $\mathbf{U}\in\operatorname{M}_{p,p}$ is an upper-triangular matrix. Then the non-linear activation function (Nair & Hinton, 2010; Huang et al., 2018) is given by
@@
Another approach for embedding graphs on Grassmann manifolds has also been proposed in Zhou et al. (2022). However, unlike our method, this method creates a Grassmann representation for a graph via a SVD of the matrix formed from node embeddings previously learned by a Euclidean neural network. Therefore, it is not designed to learn node embeddings on Grassmann manifolds.
@@ 4.1 Human Action Recognition
We use three datasets, i.e., HDM05 (Müller et al., 2007), FPHA (Garcia-Hernando et al., 2018), and NTU RBG+D 60 (NTU60) (Shahroudy et al., 2016). We compare our networks against the following state-of-the-art models: SPDNet (Huang & Gool, 2017) [^1] https://github.com/zhiwu-huang/SPDNet., SPDNetBN (Brooks et al., 2019) [^2] https://papers.nips.cc/paper/2019/hash/6e69ebbfad976d4637bb4b39de261bf7-Abstract.html., SPSD-AI (Nguyen, 2022a), GyroAI-HAUNet (Nguyen, 2022b), and MLR-AI (Nguyen & Yang, 2023).
@@ 4.2 Node Classification
We use three datasets, i.e., Airport (Zhang & Chen, 2018), Pubmed (Namata et al., 2012a), and Cora (Sen et al., 2008), each of them contains a single graph with thousands of labeled nodes. We compare our network Gr-GCN++ (see Fig. 1) against its variant Gr-GCN-ONB (see Appendix E.2.4) based on the ONB perspective. Results are shown in Tab. 2. Both networks give the best performance for $n=14$ and $p=7$. It can be seen that Gr-GCN++ outperforms Gr-GCN-ONB in all cases. The performance gaps are significant on Pubmed and Cora datasets.
@@ Appendix A Notations
~$\widetilde{\oplus}_{gr}$, ~$\widetilde{\ominus}_{gr}$ Binary and inverse operations in $\widetilde{\operatorname{Gr}}_{n,p}$
@@
.,.$\langle.,.\rangle^{ai}$ Inner product in $\operatorname{Sym}_{n}^{+,ai}$
.,.$\langle.,.\rangle^{gr}$ Inner product in $\operatorname{Gr}_{n,p}$
.,.$\langle.,.\rangle^{psd,ai}$ Inner product in $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+,ai}$
.,.$\langle.,.\rangle^{ai}_{\mathbf{P}}$ Affine-Invariant metric at $\mathbf{P}$
@@ Appendix B MLR in Structure Spaces
  /* $\mathbf{X}_{i}=\mathbf{U}_{i}\boldsymbol{\Sigma}_{i}\mathbf{V}^{T}_{i}$/
@@ 11      $\mathbf{U}^{m}\leftarrow\operatorname{GrGeodesic}(\mathbf{U}^{m},\mathbf{U},\gamma)$;
15$(\mathbf{U}^{i}_{X},\mathbf{S}^{i}_{X})_{i=1,\ldots,N}=\operatorname{GrCanonicalize}((\mathbf{X}_{i},\mathbf{U}_{i})_{i=1,\ldots,N},\mathbf{U}^{m})$;
@@ Appendix C Formulation of MLR from the Perspective of Distances to Hyperplanes
As shown in Lebanon & Lafferty (2004), Eq. (3) can be rewritten as
@@ D.1 Datasets
HDM05 (Müller et al., 2007)
It has 2337 sequences of 3D skeleton data classified into $130$ classes. Each frame contains the 3D coordinates of $31$ body joints. We use all the action classes and follow the experimental protocol in Harandi et al. (2018) in which 2 subjects are used for training and the remaining 3 subjects are used for testing.
FPHA (Garcia-Hernando et al., 2018)
It has 1175 sequences of 3D skeleton data classified into $45$ classes. Each frame contains the 3D coordinates of $21$ hand joints. We follow the experimental protocol in Garcia-Hernando et al. (2018) in which $600$ sequences are used for training and $575$ sequences are used for testing.
NTU60 (Shahroudy et al., 2016)
It has 56880 sequences of 3D skeleton data classified into $60$ classes. Each frame contains the 3D coordinates of $25$ or $50$ body joints. We use the mutual actions and follow the cross-subject experimental protocol in Shahroudy et al. (2016) in which data from 20 subjects are used for training, and those from the other 20 subjects are used for testing.
@@ D.2.1 Setup
We use the PyTorch framework to implement our networks and those from previous works. These networks are trained using cross-entropy loss and Adadelta optimizer for 2000 epochs. The learning rate is set to $10^{-3}$. The factors $\beta$ (see Proposition 3.4) and $\lambda$ (see Definition 3.9) are set to 00 and $1$, respectively. For GyroSpd++, the sizes of output matrices of the convolutional layer are set to $34\times 34$, $21\times 21$, and $11\times 1$ 11 for the experiments on HDM05, FPHA, and NTU60 datasets, respectively. For GyroSpsd++, the sizes of SPD matrices in structure spaces are set to $20\times 20$, $14\times 14$, and $8\times 8$ for the experiments on HDM05, FPHA, and NTU60 datasets, respectively. We use a batch size of 32 for HDM05 and FPHA datasets, and a batch size of 256 for NTU60 dataset.
@@ D.2.2 Input data
We use a similar method as in Nguyen (2022b) to compute the input data for our network GyroSpd++. We first identify a closest left (right) neighbor of every joint based on their distance to the hip (wrist) joint, and then combine the 3D coordinates of each joint and those of its left (right) neighbor to create a feature vector for the joint. For a given frame $t$, a mean vector $\boldsymbol{\mu}_{t}$ and a covariance matrix $\boldsymbol{\Sigma}_{t}$ are computed from the set of feature vectors of the frame and then combined to create an SPD matrix as
@@
The lower part of matrix $\log(\mathbf{Y}_{t})$ is flattened to obtain a vector $\tilde{\mathbf{v}}_{t}$. All vectors $\tilde{\mathbf{v}}_{t}$ within a time window $[t,t+c-1]$, where $c$ is determined from a temporal pyramid representation of the sequence (the number of temporal pyramids is set to 2 in our experiments), are used to compute a covariance matrix as
@@
For SPDNet and SPDNetBN, each sequence is represented by a covariance matrix (Huang & Gool, 2017; Brooks et al., 2019). The sizes of the covariance matrices are $93\times 93$, $60\times 60$, and $150\times 150$ for HDM05, FPHA, and NTU60 datasets, respectively. For SPDNet, the same architecture as the one in Huang & Gool (2017) is used with three Bimap layers. For SPDNetBN, the same architecture as the one in Brooks et al. (2019) is used with three Bimap layers. The sizes of the transformation matrices for the experiments on HDM05, FPHA, and NTU60 datasets are set to $93\times 93$, $60\times 60$, and $150\times 150$, respectively.
@@ D.2.3 Convolutional layers
$\mathbf{P}_{(i,j)}=\operatorname{concat}_{spd}(\mathbf{P}^{1}_{(i,j)},\ldots,\mathbf{P}^{L}_{(i,j)}),$

### v2

@@ 3.3 MLR in Structure Spaces
Motivated by the works in NguyenECCV22; NguyenGyroMatMans23, in this section, we aim to build MLR on SPSD manifolds. For any $\mathbf{P}\in\operatorname{S}^{+}_{n,p}$, we consider the decomposition $\mathbf{P}=\mathbf{U}_{P}\mathbf{S}_{P}\mathbf{U}^{T}_{P}$, where $\mathbf{U}_{P}\in\widetilde{\operatorname{Gr}}_{n,p}$ and $\mathbf{S}_{P}\in\operatorname{Sym}_{p}^{+}$. Each element of $\operatorname{S}^{+}_{n,p}$ can be seen as a flat $p$-dimensional ellipsoid in $\mathbb{R}^{n}$ (BonnabelSPSD13). The flat ellipsoid belongs to a $p$-dimensional subspace spanned by the columns of $\mathbf{U}_{P}$, while the $p\times p$ SPD matrix $\mathbf{S}_{P}$ defines the shape of the ellipsoid in $\operatorname{Sym}_{p}^{+}$. A canonical representation of $\mathbf{P}$ in structure space $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+}$ is computed by identifying a common subspace and then rotating $\mathbf{U}_{P}$ to this subspace. The SPD matrix $\mathbf{S}_{P}$ is rotated accordingly to reflect the changes of $\mathbf{U}_{P}$. Details of these computations are given in Appendix H.
@@
where $\widetilde{\oplus}_{gr}$ is the binary operation in $\widetilde{\operatorname{Gr}}_{n,p}$ (see Appendix G.6 for the definition of $\widetilde{\oplus}_{gr}$).
@@
where $\widetilde{\ominus}_{gr}$ is the inverse operation in $\widetilde{\operatorname{Gr}}_{n,p}$ (see Appendix G.6 for the definition of $\widetilde{\ominus}_{gr}$).
Our construction of the binary and inverse operations in $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+,g}$ is clearly advantageous compared to the method in NguyenECCV22 since this method does not preserve the information about the subspaces of the terms involved in these operations. In addition to the binary and inverse operations, we also need to define the inner product in structure spaces.
@@
where $\lambda>0$, $\langle.,.\rangle^{gr}$ is the Grassmann inner product (NguyenGyroMatMans23) (see Appendix G.7).
The key idea to generalize MLR to a Riemannian manifold is to change the margin to reflect the geometry of the considered manifold (a formulation of MLR from the perspective of distances to hyperplanes is given in Appendix C). This requires the notions of hyperplanes and margin in the considered manifold that are referred to as hypergyroplanes and pseudo-gyrodistances (NguyenGyroMatMans23), respectively. In our case, the definition of hypergyroplanes in structure spaces, suggested by Proposition 3.2, can be given below.
@@
where $\|.\|^{gr}$ and $\|.\|^{g}$ are the norms induced by the Grassmann and SPD inner products, respectively.
@@ 3.4.1 Grassmann Logarithmic Map in The Projector Perspective
The Grassmann logarithmic map is given (BATZIES201583; bendokat2020grassmann) by
@@ 3.4.2 Graph Convolutional Networks on Grassmann Manifolds
We propose to extend GCNs to Grassmann geometry using an approach similar to chami2019hyperbolic; zhao2023modeling. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, $\mathbf{x}_{i}^{l},i\in\mathcal{V}$ be the embedding of node $i$ at layer $l$ ($l=0$ indicates input node features), $\mathcal{N}(i)=\{j:(i,j)\in\mathcal{E}\}$ be the set of neighbors of $i\in\mathcal{V}$, $\mathbf{W}^{l}$ and $\mathbf{b}^{l}$ be the weight and bias for layer $l$, and $\sigma(.)$ be a non-linear activation function. A basic GCN message-passing update (zhao2023modeling) can be expressed as
@@
For the aggregation operation, the weights $w_{ij}$ can be computed using different methods (kipf2017semisupervised; william2017inductivelearning). Let $\mathbf{X}_{i}^{l}\in\operatorname{Gr}_{n,p},i\in\mathcal{V}$ be the Grassmann embedding of node $i$ at layer $l$. For feature transformation on Grassmann manifolds, we use isometry maps based on left Grassmann gyrotranslations (NguyenGyroMatMans23), i.e.,
@@
where $\mathbf{P}_{i}^{l}$ and $\mathbf{Q}_{i}^{l}$ are the input and output node features of the aggregation operation, and $k_{i,j}=|\mathcal{N}(i)|^{-\frac{1}{2}}|\mathcal{N}(j)|^{-\frac{1}{2}}$ represents the relative importance of node $j$ to node $i$. For any $\mathbf{X}\in\operatorname{Gr}_{n,p}$, let $\exp([\operatorname{Log}^{gr}_{\mathbf{I}_{n,p}}(\mathbf{X}),\mathbf{I}_{n,p}])\widetilde{\mathbf{I}}_{n,p}=\mathbf{V}\mathbf{U}$ be a QR decomposition, where $\widetilde{\mathbf{I}}_{n,p}=\begin{bmatrix}\mathbf{I}_{p}\\ 0\end{bmatrix}\in\operatorname{M}_{n,p}$, $\mathbf{V}\in\operatorname{M}_{n,p}$ is a matrix with orthonormal columns, and $\mathbf{U}\in\operatorname{M}_{p,p}$ is an upper-triangular matrix. Then the non-linear activation function (Nair2010boltzmann; HuangAAAI18) is given by
@@
Another approach for embedding graphs on Grassmann manifolds has also been proposed in zhou2022embgraphgrassmann. However, unlike our method, this method creates a Grassmann representation for a graph via a SVD of the matrix formed from node embeddings previously learned by a Euclidean neural network. Therefore, it is not designed to learn node embeddings on Grassmann manifolds.
@@ 4.1 Human Action Recognition
We use three datasets, i.e., HDM05 (hdm05), FPHA (Garcia-HernandoCVPR18), and NTU RBG+D 60 (NTU60) (Shahroudy16NTU). We compare our networks against the following state-of-the-art models: SPDNet (HuangGool17) [^1] https://github.com/zhiwu-huang/SPDNet., SPDNetBN (BrooksRieBatNorm19) [^2] https://papers.nips.cc/paper/2019/hash/6e69ebbfad976d4637bb4b39de261bf7-Abstract.html., SPSD-AI (NguyenECCV22), GyroAI-HAUNet (NguyenNeurIPS22), and MLR-AI (NguyenGyroMatMans23).
@@ 4.2 Node Classification
We use three datasets, i.e., Airport (zhang2018link), Pubmed (Namata2012QuerydrivenAS), and Cora (sen2008ccnd), each of them contains a single graph with thousands of labeled nodes. We compare our network Gr-GCN++ (see Fig. 1) against its variant Gr-GCN-ONB (see Appendix E.2.4) based on the ONB perspective. Results are shown in Tab. 2. Both networks give the best performance for $n=14$ and $p=7$. It can be seen that Gr-GCN++ outperforms Gr-GCN-ONB in all cases. The performance gaps are significant on Pubmed and Cora datasets.
@@ Appendix A Notations
$\widetilde{\oplus}_{gr}$, $\widetilde{\ominus}_{gr}$ Binary and inverse operations in $\widetilde{\operatorname{Gr}}_{n,p}$
@@
$\langle.,.\rangle^{ai}$ Inner product in $\operatorname{Sym}_{n}^{+,ai}$
$\langle.,.\rangle^{gr}$ Inner product in $\operatorname{Gr}_{n,p}$
$\langle.,.\rangle^{psd,ai}$ Inner product in $\widetilde{\operatorname{Gr}}_{n,p}\times\operatorname{Sym}_{p}^{+,ai}$
$\langle.,.\rangle^{ai}_{\mathbf{P}}$ Affine-Invariant metric at $\mathbf{P}$
@@ Appendix B MLR in Structure Spaces
/* $\mathbf{X}_{i}=\mathbf{U}_{i}\boldsymbol{\Sigma}_{i}\mathbf{V}^{T}_{i}$ */
@@ 11 $\mathbf{U}^{m}\leftarrow\operatorname{GrGeodesic}(\mathbf{U}^{m},\mathbf{U},\gamma)$;
[…]
@@ Appendix C Formulation of MLR from the Perspective of Distances to Hyperplanes
As shown in LebanonMarginClassifierICML04, Eq. (3) can be rewritten as
@@ D.1 Datasets
HDM05 (hdm05)
It has 2337 sequences of 3D skeleton data classified into $130$ classes. Each frame contains the 3D coordinates of $31$ body joints. We use all the action classes and follow the experimental protocol in Harandi2018DimensionalityRO in which 2 subjects are used for training and the remaining 3 subjects are used for testing.
FPHA (Garcia-HernandoCVPR18)
It has 1175 sequences of 3D skeleton data classified into $45$ classes. Each frame contains the 3D coordinates of $21$ hand joints. We follow the experimental protocol in Garcia-HernandoCVPR18 in which $600$ sequences are used for training and $575$ sequences are used for testing.
NTU60 (Shahroudy16NTU)
It has 56880 sequences of 3D skeleton data classified into $60$ classes. Each frame contains the 3D coordinates of $25$ or $50$ body joints. We use the mutual actions and follow the cross-subject experimental protocol in Shahroudy16NTU in which data from 20 subjects are used for training, and those from the other 20 subjects are used for testing.
@@ D.2.1 Setup
We use the PyTorch framework to implement our networks and those from previous works. These networks are trained using cross-entropy loss and Adadelta optimizer for 2000 epochs. The learning rate is set to $10^{-3}$. The factors $\beta$ (see Proposition 3.4) and $\lambda$ (see Definition 3.9) are set to 0 and $1$, respectively. For GyroSpd++, the sizes of output matrices of the convolutional layer are set to $34\times 34$, $21\times 21$, and $11\times 11$ for the experiments on HDM05, FPHA, and NTU60 datasets, respectively. For GyroSpsd++, the sizes of SPD matrices in structure spaces are set to $20\times 20$, $14\times 14$, and $8\times 8$ for the experiments on HDM05, FPHA, and NTU60 datasets, respectively. We use a batch size of 32 for HDM05 and FPHA datasets, and a batch size of 256 for NTU60 dataset.
@@ D.2.2 Input data
We use a similar method as in NguyenNeurIPS22 to compute the input data for our network GyroSpd++. We first identify a closest left (right) neighbor of every joint based on their distance to the hip (wrist) joint, and then combine the 3D coordinates of each joint and those of its left (right) neighbor to create a feature vector for the joint. For a given frame $t$, a mean vector $\boldsymbol{\mu}_{t}$ and a covariance matrix $\boldsymbol{\Sigma}_{t}$ are computed from the set of feature vectors of the frame and then combined to create an SPD matrix as
@@
The lower part of matrix $\log(\mathbf{Y}_{t})$ is flattened to obtain a vector $\tilde{\mathbf{v}}_{t}$. All vectors $\tilde{\mathbf{v}}_{t}$ within a time window [t,t+c1][t,t+c-1], where $c$ is determined from a temporal pyramid representation of the sequence (the number of temporal pyramids is set to 2 in our experiments), are used to compute a covariance matrix as
@@
For SPDNet and SPDNetBN, each sequence is represented by a covariance matrix (HuangGool17; BrooksRieBatNorm19). The sizes of the covariance matrices are $93\times 93$, $60\times 60$, and $150\times 150$ for HDM05, FPHA, and NTU60 datasets, respectively. For SPDNet, the same architecture as the one in HuangGool17 is used with three Bimap layers. For SPDNetBN, the same architecture as the one in BrooksRieBatNorm19 is used with three Bimap layers. The sizes of the transformation matrices for the experiments on HDM05, FPHA, and NTU60 datasets are set to $93\times 93$, $60\times 60$, and $150\times 150$, respectively.
@@ D.2.3 Convolutional layers
[…]